OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (xbeenchantedx):

First you can start with y=mx+b. Plug in all the variables that are already known. Since you already know the x and y coordinates as well as the slope, you can plug in the values for x, y, and m. (-1)=(4)(-3)+b Then solve for b. (-1)=(4)(-3)+b -> Multiply 4 and -3. (-1)=(-12)+b -> Isolate "b" by adding 12 to both sides of the equation to cancel out -12. 11=b Now plug in the slope and your found "b" into y=mx+b y=4x+11

OpenStudy (xbeenchantedx):

You can check this by plugging in the (x,y) values. (-1)=4(-3)+11 (-1)=(-12)+11 -1=-1